# Zhou Zhang Zhao Wang Pan Ng Zhang Ni Ma Nguyen Yip Ma Alibaba GroupSingapore RochesterUnited States of America

### v2

@@ Abstract
Current emotional text-to-speech (TTS) systems face challenges in conveying the full spectrum of human emotions, largely due to the inherent complexity of human emotions and the limited range of emotional labels in existing speech datasets. To address these limitations, this paper introduces a TTS framework that provides flexible user control over three emotional dimensions—pleasure, arousal, and dominance—enabling the synthesis of a diverse array of emotional styles. The framework leverages an emotional dimension predictor, trained solely on categorical labels from speech data and grounded in earlier psychological research, which is seamlessly integrated into a language model-based TTS system. Experimental results demonstrate that the proposed framework effectively learns emotional styles from expressive speech, eliminating the need for explicit emotion labels during TTS training, while enhancing the naturalness and diversity of synthesized emotional speech.
keywords:
Emotional text-to-speech, emotional dimension, language model
@@ 1 Introduction
However, current emotional TTS systems face challenges stemming from the complexity of human emotions and the limited range of emotional labels available in existing databases [2]. Psychological research suggests that humans can experience around 34,000 distinct emotions [3], yet current systems struggle to capture this full spectrum. Previous emotional TTS studies have primarily relied on two approaches: supervision using categorical emotion labels [4, 5] and style transfer from reference speech [6, 7]. Both approaches are constrained by the limited diversity of emotional styles present in available datasets, often resulting in synthesized speech with averaged or stereotypical emotional expressions, which restricts fine-grained control over emotion rendering [8, 9].
@@
Recent language model (LM)-based TTS frameworks [10, 11] show strong in-context learning capabilities. With sufficient training data, these models effectively render appropriate prosody for complex text inputs and exhibit robust zero-shot speaker cloning capabilities. Unlike traditional TTS task, which focus primarily on capturing a wide range of phonetic variations across speakers [12], emotional TTS emphasizes rendering diverse emotional styles, thus requiring large and varied emotional speech datasets with detailed emotion annotations. Emotional TTS systems like CosyVoice [13] and EmoCtrl-TTS [14], built on state-of-the-art language models, require thousands of hours of emotional speech data, highlighting the complexity and cost of data collection. However, publicly available emotional speech datasets are typically small [15], making them inadequate for training robust LM-based TTS system from scratch. A common strategy is to fine-tune a pre-trained TTS model using small emotional datasets [16], but this approach risks overfitting and limits generalization across emotions and speakers [17]. To address these limitations, we propose a novel framework that synthesizes more nuanced and subtle emotions not present in the dataset by controlling continuous emotional dimensions. This method offers greater flexibility for emotion rendering and leverages the in-context learning capabilities of language models, enabling the generation of a wide spectrum of human emotions, even with limited or no emotion-labeled data.
@@
In this paper, we propose an effective approach for LM-based TTS systems to generate diverse emotional styles by controlling an emotional dimension vector. Our method draws inspiration from Russell’s emotion theory [18], which identifies three independent dimensions—pleasure, arousal, and dominance—as sufficient to capture the complexity of emotional expressions across various contexts and individuals. Using these dimensions as foundational anchors, we develop a model that predicts their values. The pre-trained dimension predictor is then integrated into a LM-based TTS system to assist in next-token prediction, alongside the speaker embedding. During inference, our framework enables direct control over emotional dimensions, allowing the synthesis of diverse emotional styles without requiring emotional reference inputs. Our contributions are summarize as below:
  • We introduce a novel approach for controlling emotional dimensions in language model-based TTS systems, enabling fine-grained emotional speech synthesis;
  • Our approach learns emotional styles directly from expressive speech, eliminating the need for explicit emotion labels during TTS training;
  • During inference, the model allows users to control and explore a broad spectrum of human emotions by adjusting the emotional dimension vector. Grounded in established psychological theory, this approach ensures that the synthesized emotional styles align with the full range of human emotions;
  • We further demonstrate that our method enhances naturalness in zero-shot speaker cloning tasks, consistently outperforming the baseline system;
@@
Figure 1: An overview of the proposed text-to-speech (TTS) framework with emotional dimension control, consisting of: (a) Emotional Dimension (ED) Predictor Training, and (b) Text-to-Speech Flow. The ED predictor is pre-trained on an emotional speech dataset to map emotional features to dimension representations via anchored dimensionality reduction. It then guides the autoregressive language model (LM) to predict acoustic details. ‘P’, ‘A’, and ‘D’ denote ‘Pleasure’, ‘Arousal’, and ‘Dominance’.
@@ 2.1 Problem Formulation
We model TTS as a conditional language modeling task, where the speech waveform is quantized into discrete tokens and is then modeled with a language model [10, 13, 12]. However, training language models typically requires a large amount of data to achieve stable performance [19, 20], making it impractical as to emotional speech. Speech emotion, as a form of speech expressiveness, can be represented through continuous emotional dimensions that are learnable from expressive speech [21]. Our goal is to develop a TTS model that learns from expressive speech data during training and synthesizes speech that conveys intended emotions. Similar with that of CosyVoice [13], the overall TTS system includes an autoregressive LM decoder and a flow-matching module. The autoregressive LM decoder predicts acoustic tokens from text inputs and prompt speech using emotional dimension (ED) embeddings. Subsequently, a flow-matching module predicts the Mel-spectrum from the discrete acoustic tokens. Finally, the HiFi-GAN vocoder synthesizes the speech waveform from the Mel-spectrum.
@@ 2.2 Predicting Emotional Dimensions from Categories
Human emotions can be represented categorically [22] or dimensionally [23]. Compared to categorical emotional approach [22] that lacks the capability to capture subtle emotional differences, the dimensional approach, such as Russell’s circumplex model, views emotions as fluid experiences across several dimensions, offering a more nuanced understanding [23]. The specific definitions [18] of 151 emotion types, as rated by 300 subjects, were reliably measured on a pleasure-arousal-dominance scale as shown in Table 1. Despite those studies, most emotional TTS research still relies on a finite set of emotion labels [24, 17]. Few works have explored synthesizing mixed emotional styles by combining basic emotions [25, 26]. Current TTS systems [27, 28, 14] that control emotional dimensions rely on SER models trained with costly, labor-intensive emotional dimension annotations, and face similar biases and challenges as those encountered in SER [29, 30].
Inspired by a recent SER study [31] predicting arousal-valence only from categorical emotions, we design an ED predictor to map discrete emotions into a pleasure-dominance-arousal space (Figure 1(a)). We train the ED predictor with an emotional speech dataset, and each emotion category is assigned an anchor based on [18] to guide training, with initial vectors perturbed by a Gaussian noise $\theta$ = 0.01. Using pre-trained WavLM [32] for feature extraction, we predict emotional labels through a linear and classification layer. We derive a 128-d emotional feature vector from the linear layer and fine-tune the ED vectors from their initial settings. We create a $k$-Nearest Neighbor (kNN) graph and integrate categorical labels to differentiate emotion classes. UMAP [33] refines the vectors by minimizing cross-entropy loss between high-dimensional features and ED vectors. The ED predictor then estimates pleasure-arousal-dominance values for new speech samples, guiding acoustic predictions in the TTS model.
Table 1: Examples of the definition [18] of emotional states in terms of pleasure (‘P’), arousal (‘A’), and dominance (‘D’).
@@
Sad $-0.63$ $-0.27$ $-0.33$ Protected $0.60$ $-0.22$ $-0.40$
Surprise $0.40$ $0.67$ $-0.13$ Relaxed $0.68$ $-0.46$ $0.20$
Anxious $0.01$ $0.59$ $-0.15$ Neutral $0.00$ $0.00$ $0.00$
@@ 2.3 TTS Training with Emotional Dimension Guidance
Following CosyVoice [13], we decompose the TTS process into two stages: ‘text-to-token’ and ‘token-to-speech’. The overall TTS diagram is illustrated in Figure 1(b). Given prompt speech, emotional dimension vectors are extracted using the pre-trained ED predictor (Figure 1(a)), while the speaker embedding is obtained from a pre-trained voiceprint model [^1] https://github.com/alibaba-damo-academy/3D-Speaker/tree/main/egs/3dspeaker/sv-cam++. The input text is first processed through a Grapheme-to-Phoneme tokenizer and then mapped to a shared semantic space with speech tokens using a text encoder. These extracted features collectively serve as conditioning inputs for the speech tokenizer, which produces speech token sequences.
@@ 2.4  Emotion Rendering with Emotional Dimension Control
During inference, the emotional dimension vector can be obtained in two ways: either be predicted from the prompt speech (‘ED prediction’) or manually assigned by users (‘ED control’), allowing for flexible emotion customization. Given the input text and prompt speech, the language model (LM) autoregressively predicts a sequence of speech tokens, which encode both linguistic and prosodic information. These tokens are then processed by a pre-trained optimal-transport flow-matching model, which learns the underlying distribution of Mel spectrograms. The flow-matching model generates high-fidelity Mel spectrogram samples that are conditioned on the predicted speech tokens, ensuring smooth and natural prosody. Finally, the synthesized Mel spectrograms are passed through a HiFi-GAN vocoder, which converts them into a high-quality speech waveform. Our model is expected to generate expressive and natural speech while preserving the desired emotional characteristics, whether inferred from the prompt speech or explicitly controlled by the user.
Figure 2: Statistical analysis of pitch, energy, and spectral flux for nine synthesized emotions, obtained by modulating the emotional dimension vector within the proposed framework using ED control (’ED Control’).
@@ 3 Experiments
Table 2: Mean Opinion Score (‘MOS’) of naturalness for the zero-shot TTS task. In the proposed system, emotional dimension values are predicted from the prompt speech (‘ED Prediction’).
Ground Truth CosyVoice [13] Proposed
4.80 $\pm$ 0.08 4.36 $\pm$ 0.13 4.54 $\pm$ 0.18
Figure 3: Mean Opinion Score for assessing emotional intelligibility (‘E-MOS’). In the proposed system, ED values were predicted based on the emotional reference (‘ED Prediction’).
@@ 3.1 Experimental Setup
The emotional dimension predictor is trained on the train set from the English subset of the ESD dataset [34], which includes around 10 hours of emotional speech data across 5 basic emotions: Neutral, Angry, Happy, Sad, and Surprise. The emotional dimension predictor reported an 81% emotion classification accuracy on the ESD test set. We train the TTS framework on LibriTTS dataset [35], which consists of around 600 hours of speech data with transcriptions at a 24kHz sampling rate from 2,456 speakers. We merge the ‘train-clean’ and ‘train-other’ subsets as the training set and evaluate TTS performance on the ‘test-clean’ subset. Note that while LibriTTS only contains expressive speech, it does not include any explicit emotion labels.
To set up the experiments, for each speaker in the test set of LibriTTS, we randomly choose one speech sample from the same speaker to serve as the acoustic prompt, distinct from the speech to be synthesized, and then synthesize the speech based on the transcriptions. We compare 2 different inference settings:
  • Proposed with ED prediction, where ED values are predicted from the prompt audio;
  • Proposed with ED control, where ED values are directly derived from the literature [18] to simulate different emotions, as shown in Table 1;
All synthesized speech samples range from 3 to 10 seconds. Speech demos are publicly available online [^2] Speech Demos: https://demos46.github.io/is25/.
@@ 3.2 Implementation Details
The backbone TTS framework was implemented based on the official release of CosyVoice [13], which consists of a speech tokenizer, a text encoder, a text-to-token LM and a flow-matching module. The speech tokenizer employs a ESPNet Conformer ASR model along with a vector quantizer of a codebook size of 4096 after the first 6 encoder layers. The text encoder consists of 6 transform layers with an 8-head and 512-d attention layer. The LM contains 12 transform layers with an 8-head and 512-d attention layer. An optimal-transport conditional flow matching model (OT-CFM) [36] is employed to predict the Mel spectrograms. A HiFi-GAN [^3] https://github.com/jik876/hifi-gan vocoder recovers the speech signal from the predicted Mel spectrogram.
The emotional dimension predictor consists of a WavLM features extractor followed by a linear layer, and an anchored dimensionality reduction module. The WavLM model is fine-tuned for 100 epochs with a batch size of 64 using the Adam optimizer at a 1e-4 learning rate. Dimensionality reduction settings are a 0.1 minimum distance and 20 nearest neighbors, with a 1e-2 learning rate for point optimization.
@@ 3.3 Objective Evaluation
We compute the statistics of pitch, energy and spectral flux for all synthesized speech, as shown in Figure 2. These acoustic features play a crucial role in characterizing emotional expressiveness. Higher pitch and energy levels are commonly associated with emotions such as ‘Excitement’, where increased vocal effort and higher frequency components convey heightened arousal. Similarly, greater spectral flux, which reflects rapid spectral variations, is indicative of dynamic speech patterns, often linked to intense emotions like ‘Surprise’, which exhibit abrupt prosodic changes.
From Figure 2, we observe that emotions such as ‘Excited’, ‘Surprise’, ‘Happy’ and ‘Alert’ exhibit relatively higher values across all three acoustic features, suggesting that these emotions are characterized by greater vocal dynamism and expressiveness. In contrast, emotions like ‘Sad’, ‘Angry’, ‘Neutral’ and ‘Protected’ tend to have lower values, indicating more subdued, steady, or restrained speech patterns. For instance, ‘Sad’ is typically marked by lower pitch and energy, corresponding to a softer, more monotonous voice, while ‘Neutral’ maintains moderate values across all features, aligning with a balanced, unexpressive tone. These findings provide quantitative evidence that our model successfully synthesizes distinguishable emotions solely by adjusting the emotional dimensions. The alignment between our results and established emotional speech patterns suggests that the model effectively captures the expected emotional characterizations.
Figure 4: XAB test to assess the emotion intelligibility for 4 emotion pairs (‘Surprise’ and ‘Alert’, ‘Relaxed’ and ‘Protected’, ‘Excited’ and ‘Pleasure’, ‘Angry’ and ‘Anxious’). ED values were manually assigned based on Table 1 to simulate different emotions (‘ED Control’).

### v3

# Emotional Dimension Control in Language Model-Based Text-to-Speech: Spanning a Broad Spectrum of Human Emotions

@@ Abstract
Emotional text-to-speech (TTS) systems struggle to capture the full spectrum of human emotions due to the inherent complexity of emotional expressions and the limited coverage of existing emotion labels. To address this, we propose a language model-based TTS framework that synthesizes speech across a broad range of emotional styles. Our approach enables flexible user control along three continuous dimensions—pleasure, arousal, and dominance (PAD). To enable this, we train an emotional dimension predictor that maps categorical emotion labels in speech datasets into the PAD space, grounded in established psychological research. Importantly, while the emotional dimension predictor leverages categorical labels, the TTS framework itself does not require explicit emotion labels during training. Objective and subjective evaluations demonstrate that our framework effectively generates more expressive emotional styles and enhances both naturalness and diversity compared to baselines.
Index Terms—  Emotional text-to-speech, emotional dimension, emotion cloning, user control, language model-based text-to-speech
@@ 1 Introduction
However, current emotional TTS systems remain constrained by both the inherent complexity of human emotions and the narrow set of labels available in existing speech databases [21]. As a result, they can only generate a limited pre-defined set of emotions. Psychological research suggests that humans can experience around 34,000 distinct emotions [14], far beyond the expressive range captured by current systems. Prior research in emotional TTS has largely followed two paradigms: (1) supervised learning with categorical emotion labels [32, 25]; and (2) style transfer using reference speech [17]. Both approaches, however, suffer from the limited diversity of emotional expressions present in existing emotional speech datasets, often leading to synthesized speech that sounds average, stereotypical, and lacking in fine-grained emotional control. More recently, researchers have explored hierarchical emotion distributions for fine-grained editing and text-driven control [7, 6]. While these methods enhance fine-grained expressiveness, they remain fundamentally dependent on labeled emotion data, leaving the broader spectrum of human emotional experience underrepresented.
@@
Recent language model (LM)-based TTS frameworks [23, 27] show strong in-context learning capabilities. With sufficient training data, these models effectively render appropriate prosody for complex text inputs and exhibit robust zero-shot speaker cloning capabilities. Unlike traditional TTS tasks, which focus primarily on capturing a wide range of phonetic variations across speakers [35], emotional TTS emphasizes rendering diverse emotional styles, thus requiring large and varied emotional speech datasets with detailed emotion annotations. Emotional TTS systems like CosyVoice [4] and EmoCtrl-TTS [24], built on state-of-the-art language models, often training with thousands of hours of speech data, highlighting the cost of data collection. However, publicly available emotional speech datasets are typically small and lack of detailed annotations[31], making it difficult for training a robust LM-based TTS system from scratch. A common strategy is to fine-tune a pre-trained TTS model using small emotional datasets [29], but this approach risks overfitting and limits generalization across emotions and speakers [34]. To overcome these limitations, we propose a framework that controls speech synthesis through continuous emotional dimensions, enabling the generation of nuanced emotions beyond those present in training data. Leveraging LMs’ in-context learning, our approach enables flexible emotion rendering and broadens the set of emotions the system can synthesize, thereby expanding TTS’s expressive range even with limited or unlabeled emotional data.
@@
In this paper, we propose an effective approach for LM-based TTS systems to generate diverse emotional styles by controlling an emotional dimension vector. Our method draws inspiration from Russell’s emotion theory [15], which identifies three independent dimensions—pleasure, arousal, and dominance—as sufficient to capture the complexity of emotional expressions across various contexts and individuals. Using these dimensions as foundational anchors, we develop a model that predicts their values. The pre-trained emotion dimension (ED) predictor is then integrated into an LM-based TTS system to assist in next-token prediction, alongside the speaker embedding. During inference, our framework enables direct control over emotional dimensions, allowing the synthesis of diverse emotional styles without requiring emotional reference inputs. Our contributions are summarized as follows:
  • We propose a novel LM-based TTS framework that controls emotions through continuous emotional dimensions, moving beyond the limitations of conventional categorical labels and costly dimensional annotations. Unlike prior approaches, our model learns expressive emotional styles directly from speech, without requiring explicit emotion labels during TTS training;
  • We introduce a psychologically grounded control space, where users can directly manipulate a low-dimensional emotional vector (Pleasure–Arousal–Dominance) to explore and synthesize a wide spectrum of emotions with fine granularity, extending beyond those predefined emotions observed in training data;
  • Experiments show that our approach not only improves the naturalness of zero-shot emotion cloning but also consistently outperforms a strong LM-based TTS baseline in terms of emotional intelligibility with subjective evaluation.
@@
Fig. 1: An overview of the proposed text-to-speech (TTS) framework with emotional dimension control, consisting of: (a) Emotional Dimension (ED) Predictor Training, and (b) Text-to-Speech Flow. The ED predictor is pre-trained on an emotional speech dataset to map emotional features to dimension representations via anchored dimensionality reduction. It then guides the autoregressive language model (LM) to predict acoustic details. ‘P’, ‘A’, and ‘D’ denote ‘Pleasure’, ‘Arousal’, and ‘Dominance’. PAD values can be either inferred from prompt speech (‘Emotion Cloning’) or assigned by humans (‘Emotion Control’).
@@ 2.1 Problem Formulation
We formulate TTS as a conditional language modeling task, where the speech waveform is quantized into discrete tokens and is then modeled with a language model [23, 4, 35]. Since training language models typically requires large-scale data to achieve robust performance [9, 10], this is often impractical for emotional TTS, as emotional speech data are scarce and costly to collect. To address this, we represent speech emotion as continuous dimensions of expressiveness that can be learned directly from expressive speech data [11]. Our goal is to develop a TTS model that learns from such data during training and synthesizes speech that conveys intended emotions.
Our TTS system adopts an autoregressive language model (LM) decoder, a flow-matching module, and a HiFi-GAN vocoder. The LM decoder first generates discrete acoustic tokens from text, which the flow-matching module converts into Mel-spectrograms. Finally, HiFi-GAN synthesizes the waveform from the predicted Mel-spectrograms. Beyond this baseline design, our model introduces two key enhancements: (1) emotional dimension (ED) embeddings for fine-grained emotion control, and (2) conditioning on a reference speech to improve voice consistency.
@@ 2.2 Predicting Emotional Dimensions from Categories
Human emotions can be represented categorically [5] or dimensionally [16]. Compared to the categorical emotional approach [5] that lacks the capability to capture subtle emotional differences, the dimensional approach, such as Russell’s circumplex model, views emotions as fluid experiences across several dimensions, offering a more nuanced understanding [16]. The specific definitions [15] of 151 emotion types, as rated by 300 subjects, were reliably measured on a pleasure-arousal-dominance scale as shown in Table 1.
Despite these advances, most emotional TTS research still relies on a limited set of categorical emotion labels [1, 34]. Only a few studies have explored generating mixed emotional styles by combining basic emotions [33, 18]. Recent TTS systems that incorporate emotional dimensions [3, 8, 24] typically depend on SER models trained with costly and labor-intensive dimensional annotations. As a result, they inherit well-known SER biases—such as subjective and inconsistent labeling and imbalanced emotion distributions—which limit their ability to generalize when applied to generation tasks [22, 13].
Inspired by a recent SER study [28] predicting arousal-valence only from categorical emotions, we design an ED predictor to map discrete emotions into a pleasure-dominance-arousal space (Figure 1(a)). We train the ED predictor with an emotional speech dataset, and each emotion category is assigned an anchor based on [15] to guide training, with initial vectors perturbed by a Gaussian noise $\theta$ = 0.01. Using pre-trained WavLM [2] for feature extraction, we predict emotional labels through a linear and classification layer. We derive a 128-d emotional feature vector from the linear layer and fine-tune the ED vectors from their initial settings. We create a $k$-Nearest Neighbor (kNN) graph and integrate categorical labels to differentiate emotion classes. UMAP [12] refines the vectors by minimizing cross-entropy loss between high-dimensional features and ED vectors. The ED predictor then estimates pleasure-arousal-dominance values for new speech samples, guiding speefch token predictions in the TTS model.
Table 1: Examples of emotional state definitions in terms of the three dimensions of Pleasure (‘P’), Arousal (‘A’), and Dominance (‘D’), following [15].
@@
Sad 0.63-0.63 0.27-0.27 0.33-0.33 Protected $0.60$ 0.22-0.22 0.40-0.40
Surprise $0.40$ $0.67$ 0.13-0.13 Relaxed $0.68$ 0.46-0.46 $0.20$
Anxious $0.01$ $0.59$ 0.15-0.15 Neutral $0.00$ $0.00$ $0.00$
@@ 2.3 TTS Training with Emotional Dimension Guidance
The overall TTS diagram is illustrated in Figure 1(b). Given prompt speech, emotional dimension vectors are extracted using the pre-trained ED predictor (Figure 1(a)), while the speaker embedding is obtained from a pre-trained voiceprint model [^1] https://github.com/alibaba-damo-academy/3D-Speaker/tree/main/egs/3dspeaker/sv-cam++. The input text is first processed through a Grapheme-to-Phoneme tokenizer and then mapped to a shared semantic space with speech tokens using a text encoder. These extracted features are concatenated and serve as conditioning inputs for the speech tokenizer, which produces speech token sequences.
@@ 2.4 Emotion Rendering with Emotional Dimension Control
During inference, the emotional dimension (ED) vector can either be inferred from the prompt speech (‘emotion cloning’) or directly specified by the users (‘emotion control’), allowing fine-grained adjustment of emotional expression. Given the input text and prompt speech, the language model autoregressively generates a sequence of speech tokens that encode both linguistic and prosodic information. These tokens are then passed to a pre-trained optimal-transport flow-matching model, which maps them to the distribution of Mel spectrograms. The flow-matching model produces high-fidelity spectrograms with smooth and natural prosody, which are subsequently converted to waveforms by a vocoder. In this way, our system synthesizes expressive and natural speech while preserving the desired emotional characteristics, whether inferred from the prompt or explicitly set by the user.
Fig. 2: Statistical analysis of pitch and spectral flux for 9 emotions synthesized via ED Control in the proposed framework.
@@ 3.1 Experimental Setup
The emotional dimension predictor is trained on the training set from the English subset of the ESD dataset [30], which includes around 10 hours of emotional speech data across 5 basic emotions: neutral, angry, happy, sad, and surprise. The emotional dimension predictor reported an 81% emotion classification accuracy on the ESD test set, which is comparable to reported SER baselines on this dataset [31]. We train the TTS framework on the LibriTTS dataset [26], which consists of around 600 hours of speech data with transcriptions at a 24kHz sampling rate from 2,456 speakers. We merge the ‘train-clean’ and ‘train-other’ subsets as the training set and evaluate TTS performance on the ‘test-clean’ subset. Note that while LibriTTS only contains expressive speech, it does not include any explicit emotion labels.
For the experimental setup, we follow a prompt-based synthesis protocol. For each speaker in the LibriTTS test set, we randomly select one utterance from the same speaker as the acoustic prompt, ensuring it is distinct from the target content. The system then synthesizes speech for the test transcriptions conditioned on this prompt.
We evaluate our framework under two inference settings, which differ in how the emotional dimension (ED) values are obtained:
  • Emotion Cloning: ED values are automatically inferred from the prompt audio, enabling zero-shot emotion cloning;
  • Emotion Control: ED values are manually specified based on psychological theory [15] to synthesize speech corresponding to different emotions, as shown in Table 1;
The emotion cloning setting best demonstrates the model’s ability to generalize to unseen emotions and speakers in a zero-shot manner. The emotion control setting highlights the flexibility of the framework by allowing users to explore and manipulate emotions beyond those present in training data. The duration of each synthesized speech sample ranges from 3 to 10 seconds. Speech demos are publicly available online [^2] Speech Demos: https://demos46.github.io/emotion_pad/.
Table 2: Mean Opinion Score (‘MOS’) of naturalness for the zero-shot emotion cloning task.
Ground Truth Emotion Cloning (Proposed) Baseline CosyVoice [4]
4.80 $\pm$ 0.08 4.54 $\pm$ 0.18 4.36 $\pm$ 0.13
@@ 3.2 Implementation Details
The backbone TTS framework is similar to that of CosyVoice [4], which consists of a speech tokenizer, a text encoder, a text-to-token LM, and a flow-matching module. The speech tokenizer employs an ESPNet Conformer ASR model along with a vector quantizer of a codebook size of 4096 after the first 6 encoder layers. The text encoder consists of 6 transform layers with an 8-head and 512-d attention layer. The LM contains 12 transform layers with an 8-head and 512-d attention layer. An optimal-transport conditional flow matching model (OT-CFM) [19] is employed to predict the Mel spectrograms. A HiFi-GAN [^3] https://github.com/jik876/hifi-gan vocoder recovers the speech signal from the predicted Mel spectrogram. The emotional dimension predictor consists of a WavLM features extractor followed by a linear layer, and an anchored dimensionality reduction module. The WavLM model is fine-tuned for 100 epochs with a batch size of 64 using the Adam optimizer at a 0.0001 learning rate. Dimensionality reduction settings are a 0.1 minimum distance and 20 nearest neighbors, with a 0.01 learning rate for optimization.
@@ 3.3 Objective Evaluation
Pitch and spectral flux are well-established acoustic correlates of emotional expressiveness: higher pitch is often linked to heightened arousal in emotions such as ‘Excitement’, while greater spectral flux, reflecting rapid spectral variation, is associated with dynamic emotions like ‘Surprise’. To assess whether our system reproduces these patterns, we compute pitch and spectral flux statistics for the synthesized speech, as shown in Figure 2.
The results reveal that emotions such as ‘Excitement’, ‘Surprise’, ‘Happy’, and ‘Alert’ exhibit relatively higher values for both features, suggesting greater vocal dynamism and expressiveness. In contrast, ‘Sad’, ‘Angry’, ‘Neutral’, and ‘Protected’ show lower values, indicating more subdued or restrained speech patterns. These trends are consistent with findings in established emotion theory [15], where similar acoustic distinctions across emotions have been reported. This alignment provides evidence that our model not only generates distinguishable emotional styles but also reflects genuine acoustic correlates of human emotional expression.
Fig. 3: Mean Opinion Score for assessing emotional intelligibility (‘E-MOS’) of our proposed system for the emotion cloning task.